OpenStudy (anonymous):

From first principles, find the derivative of y=-2sin3x

OpenStudy (bradely):

f'(x) =lim (h tends zer0) (f(x+h)-f(x))/h Use this formula find derivative

OpenStudy (anonymous):

okay, i used the f(x) - f(a) to start off with but i will try this now

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

you need the "addition angle" formula for this plus you need to know a couple limits

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

ignore the \(-2\) it has no effect on the limit, you can put it in at the end then compute \[\frac{\sin(3(x+h))-\sin(3x)}{h}\] or \[\frac{\sin(3x+3h)-\sin(3x)}{h}\] using the "addition angle formula to go to \[\frac{\sin(3x)\cos(3h)+\sin(3h)\cos(3x)-\sin(3x)}{h}\]

OpenStudy (anonymous):

then break apart in to two pieces if you have a proof in your book that the derivative of sine is cosine, you can mimic that, with the only difference that this time the input is \(3x\) and not \(x\) it will work pretty much the same

OpenStudy (anonymous):

okay I'll have a look in my textbook now. thanks for you help so far!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

if I have \[\lim_{\theta \rightarrow 0} \frac{ \sin3h }{ h } = 1\] does it actually equal one? Im not sure what to do with the 3 bit...

OpenStudy (anonymous):

not it does not

OpenStudy (anonymous):

\[\lim_{h\to 0}\frac{\sin(3h)}{h}=3\] for the following reason

OpenStudy (anonymous):

oooOH i see

OpenStudy (anonymous):

\[\lim_{h\to 0}\frac{\sin(3h)}{h}=3\lim_{h\to 0}\frac{\sin(3h)}{3h}=3\times 1=3\]

OpenStudy (anonymous):

cool! thanks for explaining :)

OpenStudy (anonymous):

you did expect a three right? because you know the answer already

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

im still going, but yes i do know the answer

OpenStudy (anonymous):

YAY i got it! thanks for your help – it is greatly appreciated! love your work

OpenStudy (anonymous):

you are quite welcome i appreciate the compliment

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i got a question, if i've got \[\frac{ \cos 3h -1 }{ h }\] then as h-->0 why does cos3h--> 0?

OpenStudy (anonymous):

this is what i am referring to: